State the number of subsets of {p, q, r, s}. Explain the answer please as I don't understand how to do this.


1
Expert's answer
2021-12-30T02:51:36-0500

Answer is "2^4=16".

Explanation :

We calculate the number of subsets of the given set according to the number of elements present in the subset .

Case 1 :

The number of subsets of cardinality 0 ( cardinality means the number of elements in the given set ) = 1"=^4C_0"

Because the only subset of cardinality 0 is "\phi" .

Case 2 : the number of subsets of cardinality 1 = "^4C_1=4"

(You can also observe "\{p\} , \{q\},\{r\},\{s\}" are the only subsets of cardinality 1 of the given set ).

Case 3 : the number of subsets of cardinality 2 "=^4C_2" "=6"

( You can also observe that "\{p,q\},\{p,r\},\{p,s\},\{q,r\},\{q,s\},"

"\{r,s\}" are the only subsets of cardinality 2 of the given set.)

Case 4 : the number of subsets of cardinality 3 "=^4C_3=4"

(You can also observe that

"\{p,q,r\},\{p,q,s\},\{p,r,s\},\{q,r,s\}"

are the only subsets of cardinality 4 of the given set ).

Case 4 : the number of subsets of cardinality 4 "=^4C_4=1"

(Observe that only subset of cardinality 4 of the given set is the set itself ).

Hence the total number of subsets of the given set "=^4C_0+^4C_1+^4C_2+^4C_3+^4C_4"

"=(1+1)^4"

(Using binomial theorem )

"=2^4=16"

(You can observe that

"1+4+6+4+1=16" )